Problem:
a person can wash their car in 30 minutes. A friend can do it in 20 minutes. How long will it take them if they work together.
I really have a tough time with word problems. thank you!

Answer by stanbon(75887)

a person can wash their car in 30 minutes. A friend can do it in 20 minutes. How long will it take them if they work together.
-----------
You have to understand the difference between time and rate.
-----------
Amount of time for 1st person = 30 min./job ; rate = (1/30) job/min
That means every minute the person does 1/30th of the job.
------------------
Amount of time for the friend= 20 min/job ; rate = (1/20)job/min.
----------------
Amount of time to do job together = x min/job ; rate =(1/x) job/min
EQUATION:
rate of 1st + rate of friend = rate together
1/30 + 1/20 = 1/x
Multiply thru by 60x to get:
2x + 3x =60
5x=60
x=12 minutes (this is the time they will take working together)
Cheers,
Stan H.